### v2

# Non-autonomous iteration of polynomials in the complex plane

Abstract.
We consider a sequence $(p_{n})_{n=1}^{\infty}$ of polynomials with uniformly bounded zeros and $\deg p_{1}\geq 1$, $\deg p_{n}\geq 2$ for $n\geq 2$, satisfying certain asymptotic conditions. We prove that the function sequence $\left(\frac{1}{\deg p_{n}\cdot...\cdot\deg p_{1}}\log^{+}|p_{n}\circ...\circ p_{1}|\right)_{n=1}^{\infty}$ is uniformly convergent in $\mathbb{C}$. The non-autonomous filled Julia set $\mathcal{K}[(p_{n})_{n=1}^{\infty}]$ generated by the polynomial sequence $(p_{n})_{n=1}^{\infty}$ is defined and shown to be compact and regular with respect to the Green function. Our toy example is generated by $t_{n}=\frac{1}{2^{n-1}}T_{n},\ n\in\{1,2,...\}$, where $T_{n}$ is the classical Chebyshev polynomial of degree $n$.
2020 Mathematics Subject Classification: Primary 37F10; Secondary 30C10, 30E10, 31A15Key words and phrases: Julia sets, polynomials, Green function, Chebyshev polynomials.
1. Introduction
Behavior of iterates $p^{n}:=p^{\circ n}$ of a polynomial $p:\mathbb{C}\longrightarrow\mathbb{C}$ has been intensely studied since the nineteenth century. For an introduction to the subject see e.g. [7]. Much more recent (and much less advanced) is the study of sequences of maps of the type $p_{n}\circ...\circ p_{1}$ where each $p_{n}$, $n\in\{1,2,...\},$ is a polynomial, but not necessarily the same one. Under the name of "generalized iteration", the works [5] and [6] deal with such sequences when the coefficients of underlying polynomials satisfy some conditions (see Subsection 2.3 for more details). For a recent study of the subject see [21]. In this article we extend the dynamical study to sequences of polynomials without necessarily assuming anything about their coefficients. Instead, we work with sequences of polynomials exhibiting certain asymptotic behavior. Our class includes not only the family $\mathcal{B}$ studied in [5] and [6], but also sequences of polynomials associated with a (fixed) compact polynomially convex set in the complex plane that have been widely studied in complex approximation and interpolation theory, often in connection with logarithmic potential theory ([27], [4], [9], [2]), namely sequences satisfying the Kalmár-Walsh condition. For examples see Subsection  2.3.
Main Example.
Recall that the classical Chebyshev polynomials satisfy $T_{d_{1}}\circ T_{d_{2}}=T_{d_{2}}\circ T_{d_{1}}=T_{d_{1}d_{2}}$. All zeros of $T_{d}$ belong to the segment $[-1,1]$. The polynomial $t_{d}=\frac{1}{2^{d-1}}T_{d}$ is the monic polynomial with minimal uniform norm (the minimal polynomial) of degree $d$ on $[-1,1]$,  so
$\lim_{d\to\infty}\|t_{d}\|_{[-1,1]}^{1/d}={\rm cap}([-1,1])=\frac{1}{2}$
and
$\lim_{d\to\infty}\frac{1}{d}\log|t_{d}(z)|=g_{[-1,1]}(z)-\log 2$
locally uniformly in $\mathbb{C}\setminus[-1,1]$. Here $g_{[-1,1]}$ is the complex Green function of the segment $[-1,1]$ and ${\rm cap}([-1,1])$ is its logarithmic capacity (for the definitions and properties of these notions, see Section 2). In consequence, $\lim_{d\to\infty}\frac{1}{d}\log|T_{d}(z)|=g_{[-1,1]}(z)$ locally uniformly in $\mathbb{C}\setminus[-1,1]$.
For a sequence of integers $(d_{n})_{n=1}^{\infty}$ not smaller than 2 we get $T_{d_{n}}\circ...\circ T_{d_{1}}=T_{d_{n}...d_{1}}$, hence $(T_{d_{n}}\circ...\circ T_{d_{1}})_{n=1}^{\infty}$ is a subsequence of the sequence of all classical Chebyshev polynomials with increasing degrees and
$\lim_{n\to\infty}\frac{1}{d_{n}...d_{1}}\log\left|(T_{d_{n}}\circ...\circ T_{d_{1}})(z)\right|=g_{[-1,1]}(z)$
locally uniformly in $\mathbb{C}\setminus[-1,1]$.
By a more dynamical approach it was shown in [20] that
$\lim_{n\to\infty}\frac{1}{d_{n}...d_{1}}\log^{+}\left|(T_{d_{n}}\circ...\circ T_{d_{1}})(z)\right|=g_{[-1,1]}(z)$
and this convergence is uniform in the whole complex plane.
The function $\frac{1}{d_{n}...d_{1}}\log^{+}\left|(T_{d_{n}}\circ...\circ T_{d_{1}})\right|$ is the Green function of the preimage of the closed unit disk, $g_{(T_{d_{n}}\circ...\circ T_{d_{1}})^{-1}\left(\overline{\mathbb{D}}(0,1)\right)}$ (cf. (2.4) and Example 2.1.3), and the convergence of functions can be thought of as the convergence in Klimek’s metric of the sets $(T_{d_{n}}\circ...\circ T_{d_{1}})^{-1}\left(\overline{\mathbb{D}}(0,1)\right)$ to $[-1,1]$. It is natural to ask whether the convergence still holds if $(T_{n})_{n=1}^{\infty}$ is replaced by another sequence $(p_{n})_{n=1}^{\infty}$ of polynomials (not necessarily associated with a specific compact set $K$) and $\overline{\mathbb{D}}(0,1)$ is replaced by an arbitrary compact, polynomially convex regular set $E$. We show that this is indeed the case when the polynomials $p_{n}$ satisfy certain asymptotic condition and their zeros are uniformly bounded. Our main result is the following (see Theorem 3.4.1):
Main Theorem.
Consider a sequence of polynomials $(p_{n})_{n=1}^{\infty}$ with $\deg p_{1}\geq 1$ and $\deg p_{n}\geq 2$ for $n\geq 2$. Define $\alpha:=\limsup_{n\to\infty}\frac{1}{{\rm deg}\ p_{n}}$. Assume that there exist an $r>0$ and a function $h:\mathbb{C}\longrightarrow\mathbb{R}$ satisfying $h(z)-\alpha\log|z|\longrightarrow\infty\text{ as }|z|\to\infty$ such that
  1. $(i)$
    $\bigcup_{n=1}^{\infty}p_{n}^{-1}(0)\subset\overline{\mathbb{D}}(0,r)$;
  2. $(ii)$
    $\exists N>1\ \forall n\geq N\ \forall\zeta\in\mathbb{C}\setminus\overline{\mathbb{D}}(0,r):\;\frac{1}{\deg p_{n}}\log^{+}|p_{n}(\zeta)|>h(\zeta);$
  3. $(iii)$
    $\sup_{n\geq 1}\frac{1}{\deg p_{n}}\log^{+}|p_{n}(z)|<+\infty$ on a non-empty open set $D\subset\mathbb{C}$.
Then, for every compact, polynomially convex and regular set $E$, the sequence
$\left((p_{n}\circ...\circ p_{1})^{-1}(E)\right)_{n=1}^{\infty}$
is convergent in Klimek’s metric to the limit ${\mathcal{K}}[(p_{n})_{n=1}^{\infty}]$, independent of $E$. In particular ${\mathcal{K}}[(p_{n})_{n=1}^{\infty}]$ is non-empty, compact, regular and polynomially convex.
We call the set ${\mathcal{K}}[(p_{n})_{n=1}^{\infty}]$ the non-autonomous Julia set generated by the sequence $(p_{n})_{n=1}^{\infty}$. We establish basic properties of such Julia sets, showing that they are non-empty, compact, polynomially convex and regular. Our methods are primarily based on the approach developed in [18] (see also [19]).
We include a good amount of background material in our presentation to make it as self-contained as possible. Section 2 starts with preliminaries from logarithmic potential theory and complex approximation theory. Later, we introduce the notion of a guided sequence (such sequences satisfy conditions $(i)$ and $(ii)$ in the Main Theorem). Theorem 2.2.8 provides a characterization of guided sequences. Lemma 2.2.3 and Corollary 2.2.4 sharpen some previously available results. Examples of guided sequences appear in Subsection 2.3. Among them are Kalmár-Walsh sequences, important in approximation theory, and sequences from the class $\mathcal{B}$, previously studied in connection with generalized iteration. Using a concrete example, we point out that the class of guided sequences is larger than the class $\mathcal{B}$. More new material appears in Section 3, including the definition of a non-autonomous filled Julia set, as well as proofs of existence and properties of such sets. The approach using Klimek’s metric is explained. The presentation ends with Subsection 3.5, where we explore in more detail the case of the non-autonomous filled Julia set generated by $t_{d}=\frac{1}{2^{d-1}}T_{d},\ d\in\{1,2,...\}$, that is, by Chebyshev polynomials on the segment $[-1,1]$.
2. The Green function, the logarithmic capacity and asymptotic behavior of polynomial sequences
2.1. The Green function and related notions
Let us first define the Green function and the logarithmic capacity of a set (mostly following [25, Section II.4]). Our definitions are not the most general ones, but they still yield a notion of a regular set which agrees with the one considered in literature. In particular, finite sets are not regular.
Definition 2.1.1.
Let $D\subset\mathbb{C}$ be an unbounded domain. Consider a function $g_{D}:D\longrightarrow\mathbb{R}$ with the following properties:
  1. (1)
    $g_{D}$ is harmonic and positive in $D$;
  2. (2)
    $g_{D}(z)$ tends to 00 as $z\to\partial D$;
  3. (3)
    $g_{D}(z)-\log|z|$ tends to a finite number $\gamma$ as $z\to\infty$.
If such a function $g_{D}$ exists, it is called the Green function of $D$ with pole at infinity.
It can be proved that if the Green function exists, it is unique.
Let $K\subset\mathbb{C}$ be compact. Define the polynomially convex hull of $K$  as
$\widehat{K}:=\left\{z\in\mathbb{C}\ |\;\forall p-\mbox{ polynomial: }\;|p(z)|\leq\|p\|_{K}\right\},$
where $\|p\|_{K}:=\sup_{z\in K}|p(z)|$. Let $D_{\infty}^{K}$ denote the unbounded component of $\mathbb{C}\setminus K$. We have $\widehat{K}=\mathbb{C}\setminus D_{\infty}^{K}$. The set $K$ is called polynomially convex if $K=\widehat{K}$. Note that polynomial convexity of $K$ is equivalent to the connectedness of $\mathbb{C}\setminus K$.
Definition 2.1.2.
Let $K\subset\mathbb{C}$ be a compact set such that the Green function $g_{D_{\infty}^{K}}$ of $D_{\infty}^{K}$ with pole at infinity exists. We say then that $K$ is regular and define the Green function $g_{K}:\mathbb{C}\longrightarrow\mathbb{R}$ of $K$ via formula
$g_{K}(z)=\begin{cases}g_{D_{\infty}^{K}}(z),&\mbox{ if }z\in{D_{\infty}^{K}}\\ 0,&\text{ if }z\in\widehat{K}\end{cases}.$
The number $\exp(-\gamma)$, where $\gamma$ is the limit from Definition 2.1.1(3), is called the logarithmic capacity of $K$ and denoted by ${\rm cap}(K)$.
If $K$ is regular, then $g_{K}=g_{\widehat{K}}$ and it is a continuous function in view of Definitions 2.1.1 and 2.1.2. Moreover $g_{K}$ is subharmonic in $\mathbb{C}$.
Example 2.1.3.
([26, 2.6]) Let $a\in\mathbb{C}$, $R>0$ and let $K=\overline{\mathbb{D}}(a,R):=\{z:|z-a|\leq R\}$. Then $g_{K}(z)=\log^{+}(|z-a|/R):=\max\{0,\log(|z-a|/R)\}$ and ${\rm cap}(K)=R$.
Example 2.1.4.
([24, §1.2.b)]) $g_{[-1,1]}(z)=\log\left|z+\sqrt{z^{2}-1}\right|$, with the square root branch defined in $\mathbb{C}\setminus(-\infty,0)$ so that $\sqrt{1}=1$.
Example 2.1.5.
([24, §1.2.c)]) Fix an $R>1$. Let $K=E_{R}$ be the ellipse with foci $-1,+1$ and semiaxes $a=\frac{1}{2}\bigl{(}R+\frac{1}{R}\bigr{)},\ b=\frac{1}{2}\bigl{(}R-\frac{1}{R}\bigr{)}$. Then $g_{K}(z)=\log^{+}\left({\left|z+\sqrt{z^{2}-1}\right|}/{R}\right)$ and ${\rm cap}(K)=(a+b)/2=R/2$.
Let us note also an easy consequence of the definitions. If $K\subset\mathbb{C}$ is compact and regular, then
(2.1) $g_{K}\in\mathcal{L}^{+},$
where $\mathcal{L}^{+}$ is a special class of functions, defined below.
Definition 2.1.6 (cf. [26, Introduction]).
Let $SH(\mathbb{C})$ denote the family of all subharmonic functions on $\mathbb{C}$. We define the Lelong class
:=$\mathcal{L}:=\left\{u\in SH(\mathbb{C})\ \left|\quad\exists C=C(u)\ \forall z\in\mathbb{C}:\ u(z)\leq C+\log^{+}|z|\right.\right\}.$
An important subclass of the Lelong class is
+:=$\mathcal{L}^{+}:=\left\{u\in\mathcal{L}\ \left|\quad\exists c=c(u)\ \forall z\in\mathbb{C}:\ c+\log^{+}|z|\leq u(z)\right.\right\}.$
We will use the following result:
Proposition 2.1.7 ([26, Lemma 3.4, (4), (5) and (6)]).
Let $\mathcal{U}\subset\mathcal{L}$ be a non-empty family such that each function $v\in\mathcal{U}$ is continuous. Let $u:=\sup\{v:v\in\mathcal{U}\}$. The following are equivalent:
  1. $(i)$
    There exists a non-empty open $D\subset\mathbb{C}$ such that $u(z)<+\infty$ for all $z\in D$;
  2. $(ii)$
    $u$ is bounded from above on every compact set $K\subset\mathbb{C}$ (i.e., the family $\mathcal{U}$ is locally uniformly bounded from above).
  3. $(iii)$
    $u^{*}\in\mathcal{L}$, where $u^{*}(z):=\limsup_{\zeta\to z}u(\zeta)$.
The Green function of a compact regular set $K\subset\mathbb{C}$ can be also characterized as follows (cf. [15, the beginning of Chapter 5 and Theorem 5.1.7]) for every $z\in\mathbb{C}$:
(2.2) $g_{K}(z)=\sup\left\{u(z):\ u\in\mathcal{L},u|_{K}\leq 0\right\}=$
$=\sup\left\{\frac{1}{{\rm deg}P}\log|P(z)|:\ P\text{ a non-constant polynomial},\|P\|_{K}\leq 1\right\}$
Immediately from this characterization one can deduce that if $A,B\subset\mathbb{C}$ are compact and regular, then
(2.3) $A\subset B\quad\Longrightarrow\quad g_{A}\geq g_{B}.$
Also, the following polynomial transformation formula can be deduced from (2.2). Let $f$ be a polynomial of degree $n\geq 1$. Then
(2.4) $g_{f^{-1}(A)}=\frac{1}{n}g_{A}\circ f.$
In particular, the preimage under $f$ of a regular set, e.g., a closed disk, a line segment or an ellipse, is a regular set.
Definition 2.1.8.
Let $\varepsilon>0$ and let $K\subset\mathbb{C}$ be compact and regular. Then the $\varepsilon$-sublevel set of $g_{K}$ (also called the $\varepsilon$-augmentation of $K$) is $K_{\varepsilon}:=\left\{z\in\mathbb{C}:\;g_{K}(z)\leq\varepsilon\right\}$.
If in addition $K$ is polynomially convex, the family $\{K_{\varepsilon}\}_{\varepsilon>0}$ forms a neighbourhood base of the set $K$ in $\mathbb{C}$ as shown in [16, Corollary 1]. It was proved by M. Mazurek (published in [26, Proposition 5.11]) that
(2.5) $g_{K_{\varepsilon}}=\max(0,g_{K}-\varepsilon).$
This result implies the following properties of the sublevel sets:
Proposition 2.1.9 (cf. [3, Proposition 2.3]).
Let $K$ be a regular compact subset of $\mathbb{C}$. Then for every $\varepsilon>0$ the set $K_{\varepsilon}$ is polynomially convex and regular. Furthermore, ${\rm cap}(K_{\varepsilon})=\exp{(-\gamma+\varepsilon)}$.
2.2. Properties of sequences of polynomials
Let us start with a property of a single polynomial.
Lemma 2.2.1.
If $P:\mathbb{C}\longrightarrow\mathbb{C}$ is a polynomial and $\deg P\geq 2$, then
(2.6) $\forall\theta>1\ \exists R=R(\theta)>0:\quad|z|\geq R\quad\Longrightarrow\quad|P(z)|\geq\theta|z|.$
Proof.
Observe that ${|P(z)|}/{|z|}\longrightarrow\infty,$ if $|z|\rightarrow\infty$, since $\deg P\geq 2$. ∎
Now we turn to sequences. First we propose a modification of [8, Definition 3.1]:
Definition 2.2.2.
Let $K\subset\mathbb{C}$ be a regular compact set. Let further $(p_{n})_{n=1}^{\infty}$ be a sequence of complex polynomials such that $\deg p_{n}=n$. The sequence $(p_{n})_{n=1}^{\infty}$ is called $K$-guided if there exist $\delta>0,a>0$ and $b\in\mathbb{R}$ such that
  1. (*)
    $K~{}\cup~{}{\bigcup_{n=1}^{\infty}p_{n}^{-1}(0)}\subset\overline{\mathbb{D}}(0,\delta)$,
  2. (**)
    $\forall z\in\mathbb{C}\setminus\overline{\mathbb{D}}(0,\delta)\ \exists U$ – a neighbourhood of $z$ such that:
    $\exists N=N(U)\geq 1\ \forall n\geq N\ \forall\zeta\in U:\;\frac{1}{n}\log^{+}|p_{n}(\zeta)|\geq ag_{K}(\zeta)-b.$
Actually, in [8, Definition 3.1] $K$ belonged to a bigger family of compact sets, where a more general definition of the Green function was applicable. The authors also assumed instead of (**) that inequality
$\liminf_{n\rightarrow\infty}\frac{1}{n}\log^{+}|p_{n}(z)|\geq ag_{K}(z)-b$
should hold “locally uniformly” in $\mathbb{C}\setminus{\rm conv}K$ (instead of in $\mathbb{C}\setminus\overline{\mathbb{D}}(0,\delta)$). We prefer to write the explicit meaning of this notion, namely (**), following a personal communication with the authors.
Observe that because of the non-negativity of $g_{K}$, if $(p_{n})_{n=1}^{\infty}$ is $K$-guided with a number $a$, it is also $K$-guided with any number $a^{\prime}\in(0,a)$. In particular without loss of generality one can assume that $a\in(0,1]$. On the other hand, $a\leq 1$ necessarily for sequences satisfying an additional assumption, see Proposition 3.2.9.
Note that the assumption $\deg p_{n}=n$ can be relaxed (in particular, all polynomials may be allowed to have the same degree). The fractions $\frac{1}{n}$ in (**) and in all statements based on this condition should be then replaced by $\frac{1}{\deg p_{n}}$, $n\in\{1,2,...\}$. However, in what follows only finitely many polynomials will be allowed to be of degree 1, the others must have higher degrees. Equivalently, $\alpha:=\limsup_{n\to\infty}\frac{1}{{\rm deg}\ p_{n}}\leq 1/2$. In this case, without loss of generality we can assume that only $p_{1}$ is allowed to be of degree 1. Indeed, if $\nu_{0}:=\min\{\nu\in\mathbb{N}|\ \forall n\geq\nu:\ \deg p_{n}\geq 2\}\geq 3$, we can work instead with polynomials $P_{1}:=p_{\nu_{0}-1}\circ...\circ p_{1},P_{2}:=p_{\nu_{0}},...,P_{n}:=p_{\nu_{0}-2+n},\ n\geq 2$. Another possibility is to take $P_{1}$ to be the composition of all polynomials of degree 1 in the sequence (if any) and then put $P_{2}:=p_{k_{2}}$ with $k_{2}=\min\{n:\deg p_{n}>1\}$, $P_{3}:=p_{k_{3}}$ with $k_{3}=\min\{n\neq k_{2}:\deg p_{n}>1\}$ and so on.
We will now prove an important technical result. Note that a $K$-guided sequence (in the more general sense as explained above) satisfies the assumptions of the following Lemma 2.2.3 provided $a>\alpha$ (with a small change of $b$ if needed).
Lemma 2.2.3.
Let $(p_{n})_{n=1}^{\infty}$ be a sequence of polynomials with $\deg p_{1}\geq 1$ and $\deg p_{n}\geq 2$ for $n\in\{2,3,...\}$. Define $\alpha:=\limsup_{n\to\infty}\frac{1}{\deg p_{n}}$. Assume that there exist an $r>0$ and a function $h:\mathbb{C}\longrightarrow\mathbb{R}$ satisfying $h(z)-\alpha\log|z|\longrightarrow\infty\text{ as }|z|\to\infty$ such that
  1. $(i)$
    $\bigcup_{n=1}^{\infty}p_{n}^{-1}(0)\subset\overline{\mathbb{D}}(0,r)$;
  2. $(ii)$
    $\forall z\in\mathbb{C}\setminus\overline{\mathbb{D}}(0,r)\ \exists U$ – a neighbourhood of $z$ such that
    $\exists N=N(U)>1\ \forall n\geq N\ \forall\zeta\in U:\;\frac{1}{\deg p_{n}}\log^{+}|p_{n}(\zeta)|>h(\zeta).$
Then $\ \exists R>0\ \forall n\geq 2:\quad|z|\geq R\quad\Longrightarrow\quad|p_{n}(z)|\geq e|z|.$
Proof.
Choose $R_{0}>\max\{2,2r\}$ such that
(2.7) $h(z)-\alpha\log|z|>2\quad\text{ if }\ |z|\geq R_{0}/2.$
Note first that since $\alpha=\limsup_{n\rightarrow\infty}\frac{1}{{\rm deg}\ p_{n}}$,
(2.8) $\exists N^{\star}>1\ \forall n\geq N^{\star}:\quad\frac{1}{{\rm deg}\ p_{n}}\log R_{0}\leq\alpha\log R_{0}+1.$
By $(ii)$ if $z\in\partial\overline{\mathbb{D}}(0,R_{0})$, then there exist a neighbourhood $U$ of $z$ and an integer $N_{z}$ such that $\forall\zeta\in U\ \forall n\geq N_{z}:\;\frac{1}{{\rm deg}\ p_{n}}\log^{+}|p_{n}(\zeta)|>h(\zeta).$ By compactness of the circle there exists $N\geq N^{\star}$ such that
$\forall\zeta\in\partial\overline{\mathbb{D}}(0,R_{0})\ \forall n\geq N:\quad\frac{1}{{\rm deg}\ p_{n}}\log^{+}|p_{n}(\zeta)|>h(\zeta).$
Note that because of (2.7) the right hand side of the inequality above is greater than 2, hence we can replace $\log^{+}$ with $\log$, obtaining
(2.9) $\forall\zeta\in\partial\overline{\mathbb{D}}(0,R_{0})\ \forall n\geq N:\quad\frac{1}{{\rm deg}\ p_{n}}\log|p_{n}(\zeta)|>h(\zeta).$
Recall that $\alpha\in[0,1/2]$. Combining (2.7), (2.8) and (2.9), we get
$\forall\zeta\in\partial\overline{\mathbb{D}}(0,R_{0})\ \forall n\geq N:\quad\frac{1}{{\rm deg}\ p_{n}}\log|p_{n}(\zeta)|-\frac{1}{{\rm deg}\ p_{n}}\log|\zeta|>$
$h(\zeta)-\alpha\log|\zeta|-1>1.$
In consequence, for all $\zeta\in\partial\overline{\mathbb{D}}(0,R_{0})$
(2.10) $\forall n\geq N:\quad\log|p_{n}(\zeta)|-\log|\zeta|>1.$
Since all zeros of the polynomials $p_{n}$ are contained in $\overline{\mathbb{D}}(0,r)\varsubsetneq\overline{\mathbb{D}}(0,R_{0})$, we may apply the Minimum Principle for harmonic functions to $z\longmapsto\log|p_{n}(z)|-\log|z|$ for $n\geq N$ and conclude that (2.10) is satisfied for all $\zeta$ in the complement of the open disk $\mathbb{D}(0,R_{0})$. Thus
$\forall n\geq N:\quad|z|\geq R_{0}\quad\Longrightarrow\quad|p_{n}(z)|\geq e|z|.$
Now, by Lemma 2.2.1
$\forall j\in\{2,3,...,N-1\}\ \exists R_{j}>0:\quad|z|\geq R_{j}\quad\Longrightarrow\quad|p_{j}(z)|\geq e|z|.$
It suffices to take $R:=\max\{R_{j}:j\in\{0,2,3,...,N-1\}\}.$
Corollary 2.2.4.
If $(p_{n})_{n=1}^{\infty}$ is a sequence of polynomials satisfying the assumptions of Lemma 2.2.3, then
$\exists\varrho>0\;\forall R\geq\varrho\;\forall n\geq 2:\quad p_{n}^{-1}\left(\overline{\mathbb{D}}(0,R)\right)\subset\overline{\mathbb{D}}(0,R).$
Let us propose another definition.
Definition 2.2.5.
Consider a sequence of polynomials $(p_{n})_{n=1}^{\infty}$ such that $\deg p_{1}\geq 1$ and $\deg p_{n}\geq 2$ for $n\in\{2,3,...\}$. Define $\alpha:=\limsup_{n\to\infty}\frac{1}{\deg p_{n}}$. We say that $(p_{n})_{n=1}^{\infty}$ is a guided sequence of polynomials if there exist an $r>0$ and a function $h:\mathbb{C}\longrightarrow\mathbb{R}$ satisfying $h(z)-\alpha\log|z|\longrightarrow\infty\text{ as }|z|\to\infty$ such that
  1. $(i)$
    $\bigcup_{n=1}^{\infty}p_{n}^{-1}(0)\subset\overline{\mathbb{D}}(0,r)$;
  2. $(ii)$
    $\exists N>1\ \forall n\geq N\ \forall\zeta\in\mathbb{C}\setminus\overline{\mathbb{D}}(0,r):\;\ \frac{1}{\deg p_{n}}\log^{+}|p_{n}(\zeta)|>h(\zeta).$
Observe that if $(p_{n})_{n=1}^{\infty}$ is a guided sequence, then for every mapping $\varphi:\mathbb{N}\longrightarrow\mathbb{N}$ such that $\varphi(1)=1$, the sequence $(p_{\varphi(n)})_{n=1}^{\infty}$ is also guided. If $\deg p_{1}>1$, the assumption $\varphi(1)=1$ can be omitted.
Note that every guided sequence $(p_{n})_{n=1}^{\infty}$ satisfies the assumptions of Lemma 2.2.3. Moreover,
$\forall z\in\mathbb{C}\setminus\overline{\mathbb{D}}(0,r):\quad\liminf_{n\to\infty}\frac{1}{\deg p_{n}}\log^{+}|p_{n}(z)|\geq h(z).$
Example 2.2.6.
If $p_{n}(z):=n^{n}z^{n}$ for $n\geq 1$, then $(p_{n})_{n=1}^{\infty}$ is guided, since $\alpha=0$ and
$\forall n\geq 2\ \forall z\in\mathbb{C}\setminus\overline{\mathbb{D}}(0,1):\quad\frac{1}{n}\log^{+}|p_{n}(z)|=\log n+\log^{+}|z|>\log^{+}|z|.$
There are of course sequences that are not guided. First one can take polynomials with zeros that are not uniformly bounded, e.g. $p_{n}(z)=z^{n}-n^{n}$, $n\geq 1$. Note also the following example.
Example 2.2.7.
Let $p_{n}(z)=2^{-n^{2}}z^{n}$ for $n\geq 1$. Then for any fixed $N\geq 1$ we have $|z|\leq 2^{N}\Longrightarrow\forall n\geq N:\ |z^{n}|\leq 2^{n^{2}}$. Fix $z\in\mathbb{C}$. We have $\liminf_{n\rightarrow+\infty}\frac{1}{n}\log^{+}|p_{n}(z)|=0$, since $\frac{1}{n}\log^{+}|p_{n}(z)|\neq 0$ only for a finite number of integers $n$. Note that the zeros of $p_{n}$ are uniformly bounded.
We are ready to present the main result of this section – a characterization of guided sequences. We relax the assumptions on the degrees for the $K$-guided sequences as suggested under Definition 2.2.2.
Theorem 2.2.8.
Let $(p_{n})_{n=1}^{\infty}$ be a sequence of complex polynomials such that $\deg p_{1}\geq 1$ and $\deg p_{n}\geq 2$ for $n\in\{2,3,...\}$. Define $\alpha:=\limsup_{n\to\infty}\frac{1}{\deg p_{n}}$.
The following conditions are equivalent:
  1. $(i)$
    $(p_{n})_{n=1}^{\infty}$ is a guided sequence;
  2. $(ii)$
    $(p_{n})_{n=1}^{\infty}$ satisfies the assumptions of Lemma 2.2.3;
  3. $(iii)$
    $\exists R>1\ \forall n\geq 2:\quad p_{n}^{-1}\left(\overline{\mathbb{D}}(0,R)\right)\subset\overline{\mathbb{D}}(0,R)$;
  4. $(iv)$
    For every $K\subset\mathbb{C}$ compact and regular the sequence $(p_{n})_{n=1}^{\infty}$ is $K$-guided with $a\in(\alpha,1]$;
  5. $(v)$
    There exists $K\subset\mathbb{C}$ compact and regular such that $(p_{n})_{n=1}^{\infty}$ is $K$-guided with $a\in(\alpha,1]$.
Proof.
Note first that if $(p_{n})_{n=1}^{\infty}$ satisfies any of conditions $(i)$-$(v)$, then the zeros of the polynomials from the sequence are uniformly bounded.
$(i)\Longrightarrow(ii)$ and $(iv)\Longrightarrow(v)$ are obvious.
$(v)\Longrightarrow(ii)$ It suffices to take $h:=ag_{K}-b-1$. The condition $\lim_{|z|\to\infty}(h(z)-\alpha\log|z|)=+\infty$ holds, since $g_{K}\in\mathcal{L}^{+}$ in view of (2.1) and $a>\alpha$.
$(ii)\Longrightarrow(iii)$ This is Corollary 2.2.4.
$(iii)\Longrightarrow(iv)$ and $(iii)\Longrightarrow(i)$ Fix a compact regular set $K\subset\mathbb{C}$. Let $R>1$ be as in $(iii)$ and let $\varepsilon>0$ be such that $\overline{\mathbb{D}}(0,R)\subset K_{\varepsilon}$. We have $\bigcup_{n\geq 2}p_{n}^{-1}\left(\overline{\mathbb{D}}(0,1)\right)\subset\bigcup_{n\geq 2}p_{n}^{-1}\left(\overline{\mathbb{D}}(0,R)\right)\subset K_{\varepsilon}$.
It follows from (2.3), (2.4) and Example 2.1.3 that
$\forall z\in\mathbb{C}\ \forall n\geq 2:\quad g_{K_{\varepsilon}}(z)\leq g_{p_{n}^{-1}\left(\overline{\mathbb{D}}(0,1)\right)}(z)=\frac{1}{\deg p_{n}}\log^{+}|p_{n}(z)|.$
To get $(iv)$ with $K$ and $a\in(\alpha,1]$ it suffices now to recall that $g_{K_{\varepsilon}}=g_{K}-\varepsilon$ in $\mathbb{C}\setminus K_{\varepsilon}$ by (2.5) and that $g_{K}$ is non-negative. In view of (2.1), $(i)$ is also satisfied (e.g. with $h=g_{K_{\varepsilon}}-1$), since $\alpha\in[0,\frac{1}{2}]$. ∎
Remark 2.2.9.
A property slightly weaker than Lemma 2.2.3 and the implication $(iv)\Longrightarrow(iii)$ in Theorem 2.2.8 were proved under different assumptions as [8, Proposition 3.3].
2.3. Examples of guided sequences of polynomials
The notion of guided sequences generalizes the following classical one.
Definition 2.3.1.
Let $K\subset\mathbb{C}$ be a regular polynomially convex compact set. Consider a sequence of polynomials $(p_{n})_{n=1}^{\infty}$ such that
$p_{n}:\mathbb{C}\ni z\longmapsto\big{(}z-\zeta_{1}^{(n)}\big{)}...\big{(}z-\zeta_{n}^{(n)}\big{)}\in\mathbb{C},\quad n\in\{1,2,...\}.$
(Thus all $p_{n}$ are monic and $\forall n:\ \deg p_{n}=n.$) We say that $(p_{n})_{n=1}^{\infty}$ satisfies the Kalmár-Walsh condition, or is a KW sequence (of polynomials) associated with $K$, if the set $\bigcup_{n\in\mathbb{N}}\left\{\zeta_{1}^{(n)},...,\zeta_{n}^{(n)}\right\}$ is bounded and
(2.11) $\lim\limits_{n\to\infty}\|p_{n}\|_{K}^{1/n}={\rm cap}(K).$
When zeros of all $p_{n}$ do not have accumulation points in $\mathbb{C}\setminus K$, this is equivalent to $\lim\limits_{n\to\infty}|p_{n}(z)|^{1/n}={\rm cap}(K)\cdot\exp(g_{K}(z))$ uniformly on compact subsets of $\mathbb{C}\setminus K$ (cf. [4, Theorem 1.4], [27, Section 7.3, Theorem 3 and Section 7.4, Theorem 4] and [14, II.2.B, Theorem 1 and Lemma 1]). In particular, in this setting a KW sequence of polynomials associated with $K$ is $K$-guided, with $a=1$ and e.g. $b=-\log{\rm cap}(K)-1$. For the case when the zeros of $p_{n}$ are assumed to be uniformly bounded but are allowed to have limit points in the unbounded component of $\mathbb{C}\setminus K$ see [22] (cf. also the proof of Proposition 3.2.8).
Let us also recall the known notion of Chebyshev polynomials.
Definition 2.3.2.
Let $E\subset\mathbb{C}$ be a compact set. A monic polynomial $t_{n}$ of degree $n\geq 1$ is called the $n$th Chebyshev polynomial (or the $n$th minimal polynomial) on $E$ if $\|t_{n}\|_{E}\leq\|q\|_{E}$ for any monic polynomial $q$ of degree $n$.
An important example is associated with the interval $[-1,1]$ (cf. Main Example from the Introduction).
Example 2.3.3.
Let $E=[-1,1]$ and $n\geq 1$. Then $t_{n}:=\frac{1}{2^{n-1}}T_{n}$ is the $n$th minimal polynomial on $E$, where $T_{n}$ satisfies the formula $T_{n}\left(\frac{z+z^{-1}}{2}\right)=\frac{z^{n}+z^{-n}}{2}$. The polynomials $T_{n}$ are called the classical Chebyshev polynomials.
We observe the following fact:
Proposition 2.3.4.
If $K\subset\mathbb{C}$ is compact, polynomially convex and regular, then the sequence of Chebyshev polynomials on $K$ is a KW sequence associated with $K$.
Proof.
Fejér showed in [12] that all zeros of the Chebyshev polynomials on a compact set $K$ lie in the convex hull ${\rm conv}K$ (see also e.g. [22]). Moreover, in [13] it was proved that (2.11) holds for the sequence of minimal polynomials on a compact set. ∎
More examples of sequences $(p_{n})_{n=1}^{\infty}$ such that $\bigcup_{n\in\mathbb{N}}\left\{\zeta_{1}^{(n)},...,\zeta_{n}^{(n)}\right\}\subset K$ and (2.11) is satisfied can be found in [14, II.2].
Now we turn to other examples. The authors of [5] define a class ${\mathcal{B}}$ of sequences $(f_{n})_{n=1}^{\infty}$, where $f_{n}(z)=\sum_{j=1}^{d_{n}}a_{n,j}z^{j}$ and $d_{n}\geq 2$ for any $n$. The coefficients of polynomials in sequences in $\mathcal{B}$ satisfy some conditions, in particular
  1. (P2)
    there is a constant $A\geq 0$ such that $|a_{n,j}|\leq A|a_{n,d_{n}}|$ for $j\in\{0,...,d_{n}\}$ and all integers $n$.
It was shown in [5, 4.1] that the condition $(iii)$ in our Theorem 2.2.8 holds for sequences of class $\mathcal{B}$. Therefore they are guided. However, our methods apply also to some sequences not belonging to $\mathcal{B}$, because the condition (P2) does not hold for certain KW sequences of polynomials.
Remark 2.3.5.
The sequence of the Chebyshev polynomials on $[-1,1]$ does not belong to the class $\mathcal{B}$.
Proof.
Recall that $T_{n}$ is the classical Chebyshev polynomial of degree $n$ and the $n$th Chebyshev polynomial on $[-1,1]$ is $\frac{1}{2^{n-1}}T_{n}$ (see Example 2.3.3). We may write $T_{n}(z)=2^{n-1}z^{n}+a_{n-1}^{(n)}z^{n-1}+...+a_{0}^{(n)}$. Note that $a_{n-1}^{(n)}=0$ for every $n\geq 1$.
First we will prove that $a_{n-2}^{(n)}=-n2^{n-3}$ for $n\in\{1,2,...\}$. Indeed, this is true for $T_{1}(z)=z,\ T_{2}(z)=2z^{2}-1,\ T_{3}(z)=4z^{3}-3z$. To argue by induction, let $n$ be such that $a_{k-2}^{(k)}=-k2^{k-3}$ for $k\in\{2,3,...,n\}$. From the recurrence formula $T_{n+1}(z)=2zT_{n}(z)-T_{n-1}(z)$ (valid for all $n\geq 1$, if we set $T_{0}(z)\equiv 1$) we can express $T_{n+1}(z)$ as
$2z\left(2^{n-1}z^{n}+(-n2^{n-3})z^{n-2}+a_{n-3}^{(n)}z^{n-3}+...+a_{0}^{(n)}\right)+$
$-\left(2^{n-2}z^{n-1}+(-(n-1))2^{n-4}z^{n-3}+a_{n-4}^{(n-1)}z^{n-4}+...+a_{0}^{(n-1)}\right).$
Then $a_{n-1}^{(n+1)}=-n2^{n-2}-2^{n-2}=-(n+1)2^{n-2}$, as claimed.
Thus for the polynomial $\frac{1}{2^{n-1}}T_{n}$ the coefficient corresponding to $z^{n-2}$ is $-n/4$. Hence condition (P2) does not hold and $\left(\frac{1}{2^{n-1}}T_{n}\right)_{n=1}^{\infty}\notin\mathcal{B}.$
For the explicit expression of coefficients of $T_{n}$ corresponding to the powers of the variable $z$ see e.g. [23, T1.2 (40)].
Assume now that $(p_{n})_{n=1}^{\infty}$ is guided. Observe that it need not satisfy the Kalmár-Walsh condition, since the degrees of polynomials do not have to tend to infinity. One could ask whether the assumption $\deg p_{n}\longrightarrow\infty$ as $n\to\infty$ implies that $(p_{n})_{n=1}^{\infty}$ must be a (subsequence of a) KW sequence. In the following example, given any compact regular polynomially convex set $K\subset\mathbb{C}$, we will show that there is a guided sequence of polynomials with $\deg p_{n}\longrightarrow\infty$ as $n\to\infty$ which does not satisfy the Kalmár-Walsh condition for $K$.
Example 2.3.6.
Fix a compact regular polynomially convex set $K\subset\mathbb{C}$ and $a\in(0,1)$. We will construct a $K$-guided sequence $(P_{n})_{n=1}^{\infty}$ of polynomials satisfying estimates (**) from Definition 2.2.2 with the chosen $a$ and with $b=1$ which is not (a subsequence of) a KW sequence for $K$.
Let $L\subset\mathbb{C}$ be any compact regular polynomially convex set such that $K\cap L=\emptyset$ and $\mathbb{C}\setminus(K\cup L)$ is connected. Take $\delta>0$ such that $K\cup L\subset\overline{\mathbb{D}}(0,\delta)$. Let $(p_{n})_{n=1}^{\infty}$ and $(q_{n})_{n=1}^{\infty}$ be KW sequences of polynomials associated with $K$ and $L$ respectively, such that all zeros of $p_{n}$ lie in $K$ and all zeros of $q_{n}$ lie in $L$. Let further $k_{n},l_{n}$ be two sequences of integers such that $0<k_{n}<l_{n}$, $n\in\{1,2,...\}$, and $a=\lim_{n_{\to}\infty}k_{n}/l_{n}$. If $a=k/l\in\mathbb{Q}$ with $k,l$ coprime, simply take the sequences $k_{n},l_{n}$ to be constant, $k_{n}=k$, $l_{n}=l$. If $a\notin\mathbb{Q}$, take $l_{n}\longrightarrow\infty$ as $n\to\infty$ (e.g., by fixing a decimal expansion of $a$).
Let $P_{n}:=(p_{n})^{k_{n}}(q_{n})^{l_{n}-k_{n}}$ for $n\in\{1,2,...\}$. Then ${\rm deg}P_{n}=nl_{n}$. Furthermore,
$\lim_{n\to\infty}\frac{1}{nl_{n}}\log|P_{n}(z)|=\lim_{n\to\infty}\frac{k_{n}}{l_{n}}\frac{1}{n}\log|p_{n}(z)|+\lim_{n\to\infty}\frac{l_{n}-k_{n}}{l_{n}}\frac{1}{n}\log|q_{n}(z)|,$
which equals $ag_{K}(z)+(1-a)g_{L}(z)$ in $\mathbb{C}\setminus\overline{\mathbb{D}}(0,\delta)$ and the convergence is locally uniform. It follows that $\forall z\in\mathbb{C}\setminus\overline{\mathbb{D}}(0,\delta)\ \exists U$ – a neighbourhood of $z$ such that
$\exists N\geq 1\ \forall n\geq N\ \forall\zeta\in U:\quad\frac{1}{\deg P_{n}}\log^{+}|P_{n}(\zeta)|\geq ag_{K}(\zeta)-1.$
That is, $(P_{n})_{n=1}^{\infty}$ is a $K$-guided sequence of polynomials with the chosen value of $a$ and with $b=1$.
Suppose $ag_{K}(z)+(1-a)g_{L}(z)=g_{K}(z)$ when $z\in\mathbb{C}\setminus\overline{\mathbb{D}}(0,\delta)$. By the identity principle for harmonic functions, $ag_{K}(z)+(1-a)g_{L}(z)=g_{K}(z)$ when $z\in\mathbb{C}\setminus(K\cup L)=:\Omega$. It follows that $g_{K}\equiv g_{L}$ in $\Omega$. Thus the continuity of the Green functions and the connectedness of $\Omega$ imply that $\forall z\in\partial L:\ g_{K}(z)=0$. This is a contradiction because $K\cap L=\emptyset$.
Note that this argument also shows that $(P_{n})_{n=1}^{\infty}$ is not (a subsequence of) a KW sequence for $L$.
3. Polynomial Julia type sets
3.1. Autonomous Julia sets
Definition 3.1.1.
Let $P:\mathbb{C}\longrightarrow\mathbb{C}$ be a polynomial of degree $d\geq 2$. The (autonomous) filled Julia set of $P$ is
${\mathcal{K}}[P]:=$$\ \{z\in\mathbb{C}:\;(P^{n}(z))_{n=1}^{\infty}\text{ is bounded}\}.$
It is well known that the filled Julia set is non-empty, compact, totally invariant under $P$ (i.e., $P({\mathcal{K}}[P])=P^{-1}({\mathcal{K}}[P])={\mathcal{K}}[P]$) and
(3.1) ${\mathcal{K}}[P]:=\mathbb{C}\setminus\left\{z\in\mathbb{C}:\>\lim_{n\rightarrow\infty}P^{n}(z)=\infty\right\}.$
For more information, see e.g. [7].
Example 3.1.2.
$\forall n\geq 2:\;{\mathcal{K}}[z\longmapsto z^{n}]=\overline{\mathbb{D}}(0,1)\;\text{ and }\;{\mathcal{K}}[T_{n}]=[-1,1],$ where $T_{n}$ is the $n$th classical Chebyshev polynomial for $n\geq 2$.
Let us define the notion of escape radius of a polynomial, following [19, page 53] (a slightly different definition was proposed in [10]).
Definition 3.1.3.
An escape radius for a polynomial $P:\mathbb{C}\longrightarrow\mathbb{C}$ is a number $R>0$ with the following property
$|z|>R\quad\Longrightarrow\quad\lim_{n\rightarrow\infty}P^{n}(z)=\infty.$
Lemma 3.1.4.
$R>0$ is an escape radius for a polynomial $P$ if and only if ${\mathcal{K}}[P]\subset\overline{\mathbb{D}}(0,R).$ In particular, if $R$ is an escape radius for $P$ and $\varrho>R$, then $\varrho$ is an escape radius for $P$ too.
Proof.
This follows from the definition and from (3.1). ∎
Corollary 3.1.5.
If $R>0$ is an escape radius for a polynomial $P$, then
${\mathcal{K}}[P]=\bigcap_{n=1}^{\infty}P^{-n}\left(\overline{\mathbb{D}}(0,R)\right).$
Proof.
Let $R>0$ be such that ${\mathcal{K}}[P]\subset\overline{\mathbb{D}}(0,R)$. In consequence $\mathcal{K}[P]=P^{-n}(\mathcal{K}[P])\subset P^{-n}(\overline{\mathbb{D}}(0,R))$ for each $n\geq 1$. Conversely, if $\forall n\geq 1:\ z\in P^{-n}\left(\overline{\mathbb{D}}(0,R)\right)$, then $(P^{n}(z))_{n=1}^{\infty}\subset\overline{\mathbb{D}}(0,R)$, hence $z\in{\mathcal{K}}[P]$. ∎
A sufficient condition for a positive number to be an escape radius for a polynomial was given in Lemma 2.2.1.
3.2. Non-autonomous Julia sets
We will consider now a generalization of the autonomous Julia set. As in Definition 3.1.1, here also usually only polynomials of degree greater than 1 are considered (cf. e.g. [6], [19]). However, since we are interested in particular in KW sequences of polynomials, we will allow one polynomial to be of degree  1.
Definition 3.2.1.
Let $(p_{n})_{n=1}^{\infty}$ be a sequence of polynomials such that $\deg p_{1}\geq 1$ and $\deg p_{n}\geq 2$ for $n\geq 2$. We define the (non-autonomous) filled Julia set of the sequence $(p_{n})_{n=1}^{\infty}$ to be
${\mathcal{K}}[(p_{n})_{n=1}^{\infty}]:=$$\ \left\{z\in\mathbb{C}:\;\left((p_{n}\circ\dots\circ p_{1})(z)\right)_{n=1}^{\infty}\text{ is bounded}\right\}.$
Remark 3.2.2.
It is straightforward that in the situation from the definition
${\mathcal{K}}[(p_{n})_{n=1}^{\infty}]=p_{1}^{-1}\left({\mathcal{K}}[(p_{n})_{n=2}^{\infty}]\right).$
In particular if $p_{1}:\mathbb{C}\ni z\longmapsto z\in\mathbb{C}$, then ${\mathcal{K}}[(p_{n})_{n=1}^{\infty}]={\mathcal{K}}[(p_{n})_{n=2}^{\infty}]$.
Note that it follows from Definition 3.2.1 that
${\mathcal{K}}[(p_{n})_{n=1}^{\infty}]=\bigcup_{r\in\mathbb{N}}\bigcap_{n\geq 1}(p_{n}\circ...\circ p_{1})^{-1}\left(\overline{\mathbb{D}}(0,r)\right),$
hence ${\mathcal{K}}[(p_{n})_{n=1}^{\infty}]$ is of $F_{\sigma}$-type.
A non-autonomous filled Julia set may be finite, which is impossible for an autonomous one.
Example 3.2.3.
As shown in [6], if we take $\forall n\geq 1:\ p_{n}:\mathbb{C}\ni z\longmapsto n^{2^{n}}z^{2}\in\mathbb{C}$, then ${\mathcal{K}}[(p_{n})_{n=1}^{\infty}]=\{0\}$. Similarly, if we take $q_{1}:\mathbb{C}\ni z\longmapsto z^{2}-1\in\mathbb{C}$ and $\forall n\geq 2:q_{n}:=p_{n}$, then ${\mathcal{K}}[(q_{n})_{n=1}^{\infty}]=\{-1,1\}$. Both sequences are guided.
If $(p_{n})_{n=1}^{\infty}$ in Definition 3.2.1 above is periodic (i.e., there exists an $m$ such that $p_{m+i}=p_{i}$ for every $i$), then we obtain the autonomous filled Julia set (see Definition 3.1.1) of the polynomial $p_{m}\circ....\circ p_{1}$. In particular for a constant sequence we also get an autonomous Julia set. The interesting case is when the sequence is not periodic, e.g., when each polynomial in the sequence has a different degree.
Example 3.2.4.
If $(d_{n})_{n=1}^{\infty}$ is a sequence of integers not smaller than 2 and $p_{n}:\mathbb{C}\ni z\longmapsto z^{d_{n}}$, then ${\mathcal{K}}[(p_{n})_{n=1}^{\infty}]=\overline{\mathbb{D}}(0,1)$. Furthermore ${\mathcal{K}}[(T_{d_{n}})_{n=1}^{\infty}]=[-1,1]$. This follows from Example 3.1.2 and the composition formulae $p_{n}\circ p_{k}=p_{nk}$ and $T_{n}\circ T_{k}=T_{nk}$. In particular ${\mathcal{K}}[(z\longmapsto z^{n})_{n=1}^{\infty}]=\overline{\mathbb{D}}(0,1)$ and ${\mathcal{K}}[(T_{n})_{n=1}^{\infty}]=[-1,1]$.
Let us note a consequence of Example 3.2.4 and Remark 3.2.2.
Example 3.2.5.
For any non-constant polynomial $P:\mathbb{C}\longrightarrow\mathbb{C}$, the sets $P^{-1}\left(\overline{\mathbb{D}}(0,1\right))$ and $P^{-1}\left([-1,1]\right)$ are (non-autonomous) filled Julia sets.
The following theorem is our first result about the non-autonomous Julia set defined with the use of a guided sequence of polynomials.
Theorem 3.2.6.
Let $(p_{n})_{n=1}^{\infty}$ be a guided sequence of polynomials. Then ${\mathcal{K}}[(p_{n})_{n=1}^{\infty}]$ is non-empty and compact.
Proof.
Let $R$ be as in Lemma 2.2.3. We have
$|z|\geq R\quad\Longrightarrow\quad|(p_{n}\circ...\circ p_{2})(z)|\geq e^{n-1}|z|\longrightarrow\infty,\text{ if }n\rightarrow\infty.$
Therefore ${\mathcal{K}}[(p_{n})_{n=2}^{\infty}]=\bigcap_{n=2}^{\infty}(p_{n}\circ...\circ p_{2})^{-1}\left(\overline{\mathbb{D}}(0,R)\right)$. It follows that ${\mathcal{K}}[(p_{n})_{n=1}^{\infty}]$ is non-empty and compact by Remark 3.2.2. ∎
Recall that the Julia set obtained in this theorem may be finite, as shown in Example 3.2.3. Our aim is to find regular sets, therefore we need an additional assumption. Remark 2.3.5 shows that the approach from [5] may be insufficient in our study. Another course of action was proposed in [19], where it was required that all polynomials from the sequence $(p_{n})_{n=1}^{\infty}$ have a common escape radius $R$ such that
(3.2) $\sup_{n}\|p_{n}\|_{\overline{\mathbb{D}}(0,R)}<\infty.$
In this case ${\mathcal{K}}[(p_{n})_{n=1}^{\infty}]$ is regular. By Lemma 2.2.3, polynomials in a guided sequence have a common escape radius. But the condition (3.2) does not have to be satisfied, which can be seen in the following example.
Example 3.2.7.
Consider $p_{1}:\mathbb{C}\ni z\longmapsto z^{2}-2\in\mathbb{C}$ and $p_{n}:\mathbb{C}\ni z\longmapsto z^{n}\in\mathbb{C}$ for $n\geq 2$. The sequence $(p_{n})_{n=1}^{\infty}$ is obviously guided. Note that ${\mathcal{K}}[p_{1}]=[-2,2]$, therefore in view of Lemma 3.1.4 the smallest escape radius for $p_{1}$ is 2. Therefore $R$ is a common escape radius for $(p_{n})_{n=1}^{\infty}$ if and only if $R\geq 2$. However, $\sup_{n\geq 1}\|p_{n}\|_{\overline{\mathbb{D}}(0,2)}\geq\sup_{n}2^{n}=+\infty$. On the other hand
$\sup_{n\geq 1}\frac{1}{\deg p_{n}}\log^{+}|p_{n}(z)|=\max\left\{\frac{1}{2}\log^{+}|z^{2}-2|,\log^{+}|z|\right\}<+\infty$
on any bounded subset of $\mathbb{C}$.
We are interested in a less restrictive sufficient condition for the non-autonomous Julia set defined by a guided sequence of polynomials $(p_{n})_{n=1}^{\infty}$ to be regular. It turns out that the right condition to impose  is
(3.3) $\sup_{n\geq 1}\frac{1}{\deg p_{n}}\log^{+}|p_{n}(z)|<+\infty\text{ on an open set }D\subset\mathbb{C}.$
Obviously (3.2) implies (3.3) and Example 3.2.7 shows that the opposite implication is not true. Therefore the approach from [19] is also insufficient for our study. However, we will use some tools and results appearing there.
Before we go further let us comment more on condition (3.3). Recall that a guided sequence does not need to satisfy it, as shown in Example 2.2.6. On the other hand we have the following results. The first one confirms that the approach from [5] is insufficient for our case.
Proposition 3.2.8.
[…]
Proof.
Assume that $K\subset\mathbb{C}$ is compact, regular and polynomially convex and $(p_{n})_{n=1}^{\infty}$ is a KW sequence (see Definition 2.3.1) associated with $K$ with all zeros in $\overline{\mathbb{D}}(0,r)$ for some $r>0$. We recall a result from [22]. It was stated and proved there for Chebyshev (minimal) polynomials on a compact set $K$, but it can be easily generalized to our setting. Following the lines of the proof of that result, one can namely show that (2.11) implies locally uniform convergence on $\mathbb{C}\setminus\overline{\mathbb{D}}(0,r)$ of the sequence $(\frac{1}{n}\log|p_{n}|)_{n=1}^{\infty}$ to $g_{K}+\log{\rm cap}(K)$. This combined with the continuity of $g_{K}$ implies (3.3) for a suitable open set outside of $\overline{\mathbb{D}}(0,r)$. ∎
The second result is the promised relation for $K$-guided sequences.
Proposition 3.2.9.
[…]
Proof.
Define $u:=\sup_{n\geq 1}\frac{1}{{\rm deg}\ p_{n}}\log^{+}|p_{n}|$ and take $\delta$ from Definition 2.2.2. In view of (3.3) Proposition 2.1.7 yields $u^{*}\in\mathcal{L}$. Recall that $g_{K}\in\mathcal{L}^{+}$ by (2.1). Thus there exist $R>\delta$,  $C_{1}>0,\ C_{2}>0$ such that
$u^{*}(z)<C_{1}+\log|z|\quad\text{ and }\quad C_{2}+a\log|z|<ag_{K}(z)$
when $|z|>R$. But $u\leq u^{*}$ and $(p_{n})_{n=1}^{\infty}$ is $K$-guided, hence
$|z|>R\quad\Longrightarrow\quad C_{2}+a\log|z|-b<ag_{K}(z)-b<C_{1}+\log|z|.$
If $a>1$, take $R^{\prime}:=\max\left\{R,\exp\frac{C_{1}+b-C_{2}}{a-1}\right\}+1$ and let $|z|>R^{\prime}$. Then $(a-1)\log|z|<C_{1}+b-C_{2}$, which implies $|z|<\exp\frac{C_{1}+b-C_{2}}{a-1}<R^{\prime}$, a contradiction. ∎
3.3. Klimek’s metric
We use the following notation
${\mathcal{R}}$ $={{\mathcal{R}}}(\mathbb{C})=$
[…]
For $E,F\in{{\mathcal{R}}}$ Klimek defined in [16] their distance
(3.4) $\Gamma(E,F):=\sup_{z\in\mathbb{C}}|g_{E}(z)-g_{F}(z)|=\max\left(\sup_{z\in E}g_{F}(z),\sup_{z\in F}g_{E}(z)\right)$
and showed that $({{\mathcal{R}}},\Gamma)$ is a complete metric space. Note that a sequence $(E_{n})_{n=1}^{\infty}$ is convergent to $F$ in $({{\mathcal{R}}},\Gamma)$ if and only if $g_{E_{n}}\rightrightarrows g_{F}$, i.e. the function sequence $(g_{E_{n}})_{n=1}^{\infty}$ is uniformly convergent to $g_{F}$ in the whole complex plane.
Fix now a polynomial $P$ of degree $d\geq 1$ and consider the mapping
(3.5) $A_{P}:{\mathcal{R}}\ni K\longmapsto P^{-1}(K)\in{\mathcal{R}}.$
(2.4) yields that $A_{P}$ is an isometry if $d=1$ (since $P$ is bijective) and a contraction with contraction ratio $1/d$ if $d\geq 2$. In the latter case, since $({\mathcal{R}},\Gamma)$ is a complete metric space, by Banach Contraction Principle $A_{P}$ has a unique fixed point. This fixed point is the above defined (Definition 3.1.1) filled Julia set ${\mathcal{K}}[P]$ (see [16]). In particular ${\mathcal{K}}[P]\in{\mathcal{R}}$. Moreover, by the classical proof of Banach Contraction  Principle
$\forall E\in{\mathcal{R}}:\quad\lim_{n\rightarrow\infty}P^{-n}(E)=\lim_{n\rightarrow\infty}(A_{P})^{n}(E)={\mathcal{K}}[P].$
Once again using (2.4) we deduce that $\forall E\in{\mathcal{R}}:\;\frac{1}{d^{n}}g_{E}\circ P^{n}\rightrightarrows g_{{\mathcal{K}}[P]}.$
Some properties of the filled Julia set of a polynomial of degree at least 2 (see Definition 3.1.1) followed from the Banach Contraction Principle. We will use a generalization of this result.
Theorem 3.3.1 (Enhanced version of Banach’s Contraction Principle, [18, Lemma 4.5]).
Let $(X,\rho)$ be a complete metric space and let $(H_{n})_{n=1}^{\infty}$ be a sequence of contractions of X with contraction ratios not greater than $L<1$. If   $\forall x\in X:\;\sup_{n\geq 1}\rho(H_{n}(x),x)<\infty,$ then there exists a unique point $c\in X$ such that the sequence $(H_{1}\circ...\circ H_{n})_{n=1}^{\infty}$ converges pointwise to $c$.
Let us also quote the following result.
Proposition 3.3.2 ([17, Proposition 1]).
Let $P_{n}:\mathbb{C}\longrightarrow\mathbb{C}$ be a polynomial of degree $d_{n}\geq 2$ for $n\in\{1,2,...\}$. Let $E\in{\mathcal{R}}$ and define $E_{n}:=(P_{n}\circ...\circ P_{1})^{-1}(E)$ for $n\in\{1,2,...\}$. If
(3.6) $\sum_{n=1}^{\infty}\frac{\Gamma(P_{n+1}^{-1}(E),E)}{d_{1}d_{2}\cdots d_{n}}<\infty,$
[…]
$F=\bigcap_{n\geq 1}E_{n}=\{z\in E:\;(P_{n}\circ...\circ P_{1})(z)\in E\text{ for all }n\geq 1\}.$
Now we would like to use the Klimek metric in our case of guided sequences of polynomials. The following proposition is an important step in proving that the compact set ${\mathcal{K}}[(p_{n})_{n=1}^{\infty}]$ obtained in Theorem 3.2.6 is regular and polynomially convex. We assume additionally that (3.3) holds.
Proposition 3.3.3.
[…]
$\exists C>0\ \forall n\geq 1:\quad\Gamma(E,p_{n}^{-1}(E))\leq C.$
Proof.
Fix $E\in{\mathcal{R}}$ and $\varepsilon>\max\{0,\log{\rm cap}(E)\}$. Take $\varrho$ from Corollary 2.2.4, fix $R\geq\varrho$ big enough to satisfy $E\subset\overline{\mathbb{D}}(0,R)$ and such that
(3.7) $|z|>R\quad\Longrightarrow\quad g_{E}(z)\leq\log|z|-\log{\rm cap}(E)+\varepsilon$
(this is possible in view of Definitions 2.1.1 and 2.1.2). By Corollary 2.2.4
$p_{n}^{-1}(E)\subset p_{n}^{-1}\left(\overline{\mathbb{D}}(0,R)\right)\subset\overline{\mathbb{D}}(0,R)\quad\text{ for }n\geq 2.$
Hence
(3.8) $\forall n\geq 2:\qquad\sup_{z\in p_{n}^{-1}(E)}g_{E}(z)\leq\sup_{z\in\overline{\mathbb{D}}(0,R)}g_{E}(z)=:C_{1}.$
Note that $C_{1}$ is a non-negative number and does not depend on $n$.
The sequence $(p_{n})_{n=1}^{\infty}$ satisfies (3.3), hence by Proposition 2.1.7,
$\exists C_{2}\ \forall z\in E\ \forall n\geq 1:\quad\frac{1}{\deg p_{n}}\log|p_{n}(z)|\leq\frac{1}{\deg p_{n}}\log^{+}|p_{n}(z)|\leq C_{2}.$
Therefore $\forall n\geq 1\ \forall z\in E:\;|p_{n}(z)|\leq C_{3}^{\deg p_{n}},$ where $C_{3}=\exp(C_{2})>1$. In view of (2.4) we have
$\forall n\geq 1:\quad$ $\sup_{z\in E}g_{p_{n}^{-1}(E)}(z)=\sup_{z\in E}\frac{1}{\deg p_{n}}g_{E}(p_{n}(z))\leq$
(3.9) $\leq\sup_{z\in\overline{\mathbb{D}}(0,C_{3}^{\deg p_{n}})}\frac{1}{\deg p_{n}}g_{E}(z)=\sup_{z\in\partial\overline{\mathbb{D}}(0,C_{3}^{\deg p_{n}})}\frac{1}{\deg p_{n}}g_{E}(z),$
and the last equality follows from the Maximum Principle.
Consider the set
$\mathcal{N}:=\{n\in\mathbb{N}:\;C_{3}^{\deg p_{n}}\leq R\}.$
By the non-negativity of $g_{E}$ and the definition of $C_{1}$ in (3.8)
$\forall n\in\mathcal{N}\ \forall\zeta\in\partial\overline{\mathbb{D}}(0,C_{3}^{\deg p_{n}}):\quad\frac{1}{\deg p_{n}}g_{E}(\zeta)\leq g_{E}(\zeta)\leq C_{1}.$
Combining this with (3.9) gives
(3.10) $\forall n\in\mathcal{N}:\quad\sup_{z\in E}g_{p_{n}^{-1}(E)}(z)\leq C_{1}.$
On the other hand, (3.7) implies
$\forall n\notin\mathcal{N}\ \forall\zeta\in\partial\overline{\mathbb{D}}(0,C_{3}^{\deg p_{n}}):\quad g_{E}(\zeta)\leq\deg p_{n}\log C_{3}-\log{\rm cap}(E)+\varepsilon.$
Combining this with (3.9) gives
(3.11) $\forall n\notin\mathcal{N}:\quad\sup_{z\in E}g_{p_{n}^{-1}(E)}(z)\leq C_{4}:=C_{2}-\log{\rm cap}(E)+\varepsilon.$
Note that $C_{4}>0$ by the choice of $\varepsilon$.
It follows from (3.8), (3.10) and (3.11) that $\forall n\geq 1:\;\Gamma(E,p_{n}^{-1}(E))\leq\max\{C_{1},C_{4}\}$. ∎
3.4. Julia sets for guided sequences
We will now apply the results from Subsection 3.3 to a sequence of contractions defined in (3.5).
Theorem 3.4.1.
[…]
$\left(A_{p_{1}}\circ...\circ A_{p_{n}}\right)_{n=1}^{\infty}$
converges pointwise in $(\mathcal{R},\Gamma)$ to a constant mapping with the value ${\mathcal{K}}[(p_{n})_{n=1}^{\infty}]$. In particular this limit set is polynomially convex and regular.
Proof.
By Theorem 3.3.1 and Proposition 3.3.3, for every $E\in\mathcal{R}$ the sequence $\left(A_{p_{2}}\circ...\circ A_{p_{n}}(E)\right)_{n=2}^{\infty}$ is convergent to the same set $F\in{\mathcal{R}}$.
Take $\varrho>0$ from Corollary 2.2.4. Proposition 3.3.2 yields
$\forall R\geq\varrho:\quad F=\bigcap_{k=1}^{\infty}(p_{n}\circ...\circ p_{2})^{-1}\left(\overline{\mathbb{D}}(0,R)\right).$
Hence $F={\mathcal{K}}[(p_{n})_{n=2}^{\infty}]$ by Theorem 3.2.6.
If $\deg p_{1}>1$, in the same way (renumbering the sequence) we obtain that $\left(A_{p_{1}}\circ...\circ A_{p_{n}}(E)\right)_{n=1}^{\infty}$ is convergent to ${\mathcal{K}}[(p_{n})_{n=1}^{\infty}]$.
If $\deg p_{1}=1$, the mapping $A_{p_{1}}$ is an isometry, hence the convergence of $\left(A_{p_{2}}\circ...\circ A_{p_{n}}(E)\right)_{n=2}^{\infty}$ to ${\mathcal{K}}[(p_{n})_{n=2}^{\infty}]$ implies the convergence of $\left(A_{p_{1}}\circ...\circ A_{p_{n}}(E)\right)_{n=1}^{\infty}$ to $p_{1}^{-1}({\mathcal{K}}[(p_{n})_{n=2}^{\infty}])={\mathcal{K}}[(p_{n})_{n=1}^{\infty}]$, this last equality follows from Remark 3.2.2. ∎
Corollary 3.4.2.
[…]
In particular the function sequence
$\left(\frac{1}{\deg p_{n}\cdot...\cdot\deg p_{1}}\log^{+}\left|p_{n}\circ...\circ p_{1}\right|\right)_{n=1}^{\infty}$
is uniformly convergent in $\mathbb{C}$.
Proof.
It follows directly from Theorem 3.4.1 and the definition of Klimek’s metric (3.4). The last assertion is a consequence of (2.4) and the formula for the Green function of the unit disk (cf. Example 2.1.3). ∎
The following approximation of the non-autonomous Julia set by the autonomous Julia sets of compositions can be easily shown (cf. [1, Proposition 5]).
Corollary 3.4.3.
[…]
Remark 3.4.4.
Assumption (3.3) cannot be dropped in the results above. The sequence $(p_{n})_{n=1}^{\infty}$ from Example 2.2.6 does not satisfy this condition and the non-autonomous Julia set associated with it is $\{0\}\notin\mathcal{R}$.
We recall now another result due to Klimek.
Theorem 3.4.5 ([16, Corollary 5]).
$\forall E,F\in{\mathcal{R}}:\quad\left|\log{\rm cap}(E)-\log{\rm cap}(F)\right|\leq\Gamma(E,F).$
In particular the logarithmic capacity is continuous on $({\mathcal{R}},\Gamma)$.
Recall that ${\rm cap}\left(\overline{\mathbb{D}}(0,1)\right)=1$, moreover because of (2.4) we have
(3.12) ${\rm cap}\left(f^{-1}\left(\overline{\mathbb{D}}(0,1)\right)\right)=1$
for any non-constant monic polynomial $f$ too.
Corollary 3.4.6.
[…]
Proof.
In view of Theorem 3.4.1,
$\Gamma\big{(}\left(p_{n}\circ...\circ p_{1}\right)^{-1}\left(\overline{\mathbb{D}}(0,1)\right),{\mathcal{K}}\left[(p_{n})_{n=1}^{\infty}\right]\big{)}\longrightarrow 0\quad\text{ as }\quad n\rightarrow\infty.$
The assertion follows from (3.12) and Theorem 3.4.5. ∎
3.5. The case of Chebyshev polynomials on $[-1,1]$ (cf. Main Example)
Example 3.5.1.
The following pictures (prepared by Maciej Klimek) show approximations of ${\mathcal{K}}[(t_{n})_{n=1}^{\infty}]$, where $(t_{n})_{n=1}^{\infty}$ is the sequence of minimal polynomials on $[-1,1]$ (see Example 2.3.3). The sets depicted here are, from left to right:
  • $(t_{8}\circ...\circ t_{2}\circ t_{1})^{-1}([-1,1]\times[-0.0005,0.0005])$, (which is used as an approximation of $(t_{8}\circ...\circ t_{2}\circ t_{1})^{-1}([-1,1])$),
  • $(t_{5}\circ...\circ t_{2}\circ t_{1})^{-1}\left(\overline{\mathbb{D}}(0,1)\right)$,
  • $(t_{100}\circ...\circ t_{2}\circ t_{1})^{-1}\left(\overline{\mathbb{D}}(0,1)\right)$.
[Uncaptioned image][Uncaptioned image][Uncaptioned image]
Observe some simple geometric properties of the set ${\mathcal{K}}[(t_{n})_{n=1}^{\infty}]$.
Property 1: $[-1,1]\subset{\mathcal{K}}[(t_{n})_{n=1}^{\infty}]$. In particular, ${\mathcal{K}}[(t_{n})_{n=1}^{\infty}]\cap\mathbb{R}\neq\emptyset$.
Proof.
The set $I=[-1,1]$ is totally invariant under every polynomial $T_{n}$. Hence $t_{1}(I)=T_{1}(I)=I$, $t_{2}(I)=(1/2)T_{2}(I)=[-1/2,1/2]\subset I$, $t_{n}(I)=(1/2^{n-1})T_{n}(I)=[-1/2^{n-1},1/2^{n-1}]\subset I$, and $(t_{n}\circ...\circ t_{2}\circ t_{1})(I)\subset I$ for every $n\geq 1$. ∎
Property 2: If $z\in{\mathcal{K}}[(t_{n})_{n=1}^{\infty}]$, then $\{-z,\bar{z}\}\subset{\mathcal{K}}[(t_{n})_{n=1}^{\infty}]$.
Proof.
We have $t_{1}(z)=z$ and $t_{2}(-z)=t_{2}(z)$ for every $z\in\mathbb{C}$. Hence $(t_{n}\circ...\circ t_{2}\circ t_{1})(-z)=(t_{n}\circ...\circ t_{2}\circ t_{1})(z)$ for every $n\geq 2$ and every $z\in\mathbb{C}$. Moreover, all $t_{n}$ have real coefficients, so $(t_{n}\circ...\circ t_{1})(\overline{z})=\overline{(t_{n}\circ...\circ t_{1})(z)}$ for every $n\geq 2$ and every $z\in\mathbb{C}$. It follows that each of the sequences $\left((t_{n}\circ...\circ t_{1})(\overline{z})\right)_{n=1}^{\infty}$ and $\left((t_{n}\circ...\circ t_{1})(-z)\right)_{n=1}^{\infty}$ is bounded if and only if $\left((t_{n}\circ...\circ t_{1})(z)\right)_{n=1}^{\infty}$ is. ∎
Property 3: $\exists R\geq 2:\ {\mathcal{K}}[(t_{n})_{n=1}^{\infty}]\subset E_{R}$, where $E_{R}$ is the filled ellipse with foci ±$\pm 1$ and semiaxes $a_{R}=\frac{1}{2}(R+\frac{1}{R}),\ b_{R}=\frac{1}{2}(R-\frac{1}{R})$.
Proof.
When $R>1$, such (filled) ellipses are sublevel sets (see Definition 2.1.8 and Examples 2.1.4 and 2.1.5) of the Green function $g_{I}$, which tends to infinity as $|z|\to\infty$. Hence we have $\mathbb{C}=\bigcup_{R>1}E_{R}$, and, by compactness, $\exists R>1:{\mathcal{K}}[(t_{n})_{n=2}^{\infty}]\subset E_{R}$. The capacity of $E_{R}$ is $(a_{R}+b_{R})/2=R/2$ (see again Example 2.1.5). By monotonicity of capacity and Corollary 3.4.6, we need $R\geq 2$ for the inclusion ${\mathcal{K}}[(t_{n})_{n=2}^{\infty}]\subset E_{R}$. ∎
Property 4: ${\mathcal{K}}[(t_{n})_{n=1}^{\infty}]\subset\!\!\!\!\!/\ E_{2}$.
Proof.
Note that $E_{2}$ has the major semiaxis $a_{2}=5/4$ and the minor semiaxis $b_{2}=3/4$. We need to find a point in ${\mathcal{K}}[(t_{n})_{n=1}^{\infty}]\setminus E_{2}$.
Let us start with the following facts for $n\in\{1,2,...\}$:
  1. (a)
    $\forall x\in\mathbb{R}:\;t_{n}(-x)=(-1)^{n}t_{n}(x)$;
  2. (b)
    $t_{n}$ is increasing in the interval $(1,+\infty)$;
  3. (c)
    $\max_{z\in E_{2}}|t_{n}(z)|=|t_{n}(5/4)|=|t_{n}(-5/4)|=1+2^{-2n}\leq 5/4$.
(a) and (b) are known. To prove (c), we first compute $\max_{w\in E_{2}}|T_{n}(w)|$ (cf. [11], [14]). Recall that the classical Chebyshev polynomials satisfy the relation $T_{n}\left(\frac{z+z^{-1}}{2}\right)=\frac{z^{n}+z^{-n}}{2}$ for $n\in\{1,2,...\}.$ For $z=2e^{i\theta}$ with $\theta\in[0,2\pi)$ we thus have
$T_{n}\left(\frac{z+z^{-1}}{2}\right)$ $=\frac{2^{n}e^{in\theta}+2^{-n}e^{-in\theta}}{2}$
$=\frac{1}{2}\left((2^{n}+2^{-n})\cos n\theta+i(2^{n}-2^{-n})\sin n\theta\right).$
Then
$\left|T_{n}\left(\frac{z+z^{-1}}{2}\right)\right|^{2}=\frac{1}{4}\left(2^{2n}+2\cos 2n\theta+2^{-2n}\right)$
achieves its maximum when $\theta=0$ or $\theta=\pi$. Checking values for the corresponding $z=2$ or $z=-2$ we get
$\max_{w\in E_{2}}\left|T_{n}(w)\right|=\left|T_{n}\left(\frac{2+2^{-1}}{2}\right)\right|=\left|T_{n}\left(-\frac{2+2^{-1}}{2}\right)\right|=2^{n-1}+2^{-(n-1)}.$
Hence (c) is proved.
Observe now that Property 1 together with (a), (b) and (c) implies that $[-5/4,5/4]\subset{\mathcal{K}}[(t_{n})_{n=1}^{\infty}]$. Indeed, for every $n\geq 1$ we have
$t_{n}\left([-5/4,5/4])=t_{n}([-5/4,-1]\cup[-1,1]\cup[1,5/4]\right)\subset[-5/4,5/4],$
consequently $(t_{n}\circ...\circ t_{1})([-5/4,5/4])\subset[-5/4,5/4]$ and we get the inclusion $[-5/4,5/4]\subset{\mathcal{K}}[(t_{n})_{n=1}^{\infty}]$.
Consider the point $z_{0}=4i/5$, which does not belong to $E_{2}$ (since the minor semiaxis of $E_{2}$ is $b_{2}=3/4<4/5$). Now,
$t_{2}(z_{0})=z_{0}^{2}-1/2=-57/50\in(-5/4,-1),$
hence, for every $n\geq 2$ we have $(t_{n}\circ...\circ t_{2}\circ t_{1})(z_{0})\in[-5/4,5/4]$, and so $z_{0}\in{\mathcal{K}}[(t_{n})_{n=1}^{\infty}]$. ∎
Acknowledgements.
Both authors thankfully acknowledge their participation in Thematic Research Programme “Modern holomorphic dynamics and related fields”, Excellence Initiative – Research University programme at the University of Warsaw (a mini-semester in spring 2023). The paper was partially written thanks to the support from the programme.
The second named author extends her thanks to the Faculty of Mathematics, Informatics and Mechanics of the University of Warsaw for supporting her participation in the thematic semester “Dynamical Systems. Topological, smooth and holomorphic dynamics, ergodic theory, fractals” in Stefan Banach International Mathematical Center at the Institute of Mathematics of the Polish Academy of Sciences in Warsaw (part of “Simons Semesters in Banach Center: 2020s vision.”) and in the conference “Complex dynamics: connections to other fields”, at the University of Warsaw Conference Center in Chęciny, Poland, as well as to the Chair of Approximation, Institute of Mathematics, Jagiellonian University, Kraków, for hosting her in March – June 2023 during her study leave from the American Mathematical Society.
We would also like to thank Christian Henriksen for the discusion on the notion of $K$-guided sequences and Maciej Klimek for preparing the figures for us.